# Understanding Representation Dynamics of Diffusion Models via Low-Dimensional Modeling

### v2

@@ Abstract
Diffusion models, though originally designed for generative tasks, have demonstrated impressive self-supervised representation learning capabilities. A particularly intriguing phenomenon in these models is the emergence of unimodal representation dynamics, where the quality of learned features peaks at an intermediate noise level. In this work, we conduct a comprehensive theoretical and empirical investigation of this phenomenon. Leveraging the inherent low-dimensionality structure of image data, we theoretically demonstrate that the unimodal dynamic emerges when the diffusion model successfully captures the underlying data distribution. The unimodality arises from an interplay between denoising strength and class confidence across noise scales. Empirically, we further show that, in classification tasks, the presence of unimodal dynamics reliably indicates generalization: it emerges when the model generalizes and gradually transitions to a monotonically decreasing curve as the model begins to memorize the training data.
@@ 1 Introduction
Diffusion models, a class of score-based generative models, have achieved great empirical success in various tasks such as image and video generation, speech and audio synthesis, and solving inverse problems (Alkhouri et al., 2024; Ho et al., 2020; Rombach et al., 2022; Zhang et al., 2024; Bar-Tal et al., 2024; Ho et al., 2022; Kong et al., 2020, 2021; Roich et al., 2022; Ruiz et al., 2023; Chen et al., 2024a; Chung et al., 2022; Song et al., 2024; Li et al., 2024b). In general, these models, consisting of forward and backward processes, learn data distributions by simulating the non-equilibrium thermodynamic diffusion process (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021). Specifically, the forward process progressively adds Gaussian noise to training samples until they are fully corrupted, while the backward process involves learning a score-based model to generate samples from the noisy inputs (Hyvärinen and Dayan, 2005; Song et al., 2021).
Beyond their strong generative capabilities, recent studies have revealed that diffusion models also possess remarkable representation learning capabilities (Baranchuk et al., 2022; Xiang et al., 2023; Mukhopadhyay et al., 2023; Chen et al., 2024b; Tang et al., 2023; Han et al., 2024). In particular, the internal feature extractors of trained diffusion models have been shown to serve as powerful self-supervised learners, achieving strong performance on downstream tasks such as classification (Xiang et al., 2023; Mukhopadhyay et al., 2023), semantic segmentation (Baranchuk et al., 2022), and image alignment (Tang et al., 2023). In many cases, these representations match or even surpass those from specialized self-supervised methods, implying that diffusion models have the potential to unify generative and discriminative learning in vision.
Alongside the empirical success, a prevalent phenomenon, which we refer to as the unimodal representation dynamics, has been widely observed in diffusion models for representation learning (Baranchuk et al., 2022; Xiang et al., 2023; Tang et al., 2023): the quality of learned representations, as measured by downstream task performance, follows a unimodal trend across noise levels. Specifically, the most effective features consistently appear at an intermediate noise level, while performance degrades as inputs become either fully noisy or entirely clean (see Figure 1). Despite being widely observed, the underlying cause of this unimodal representation dynamic remains poorly understood.
@@
Figure 1: Unimodal representation dynamics in diffusion-based representation learning tasks. This unimodal representation pattern has been previously observed in diffusion-based representation learning tasks; see (Baranchuk et al., 2022; Xiang et al., 2023; Tang et al., 2023). To verify this, we train diffusion models on various datasets and evaluate downstream performance using noisy images $\bm{x}_{t}$ at different timesteps $t$. In both classification and segmentation tasks, the performance consistently follows a unimodal trend, peaking at intermediate noise levels. In (b), "mIoU" denotes mean Intersection over Union, a standard metric used in segmentation tasks.
@@
In this work, we investigate the emergence of unimodal representation dynamics in diffusion models and its cause. Our analysis characterizes how representation quality varies across noise levels and offers new perspectives on how diffusion models, despite their generative design, can excel at representation learning. Specifically, we conduct a comprehensive theoretical and empirical study of unimodal representation dynamics in diffusion models. Motivated by the well-established observation that natural image data typically reside on a low-dimensional manifold (Gong et al., 2019; Pope et al., 2021; Stanczuk et al., 2022), we examine the representation dynamics through how diffusion models learn from noisy mixtures of low-rank Gaussian (MoLRG) distributions. We show that the emergence of unimodal representations is intrinsically linked to the model’s ability to capture the underlying low-dimensional structure of the data. Our key contributions are as follows:
@@
Recent empirical advances in leveraging diffusion models for downstream representation learning have gained significant attention. However, a theoretical understanding of how diffusion models learn representations across different noise levels remains largely unexplored. Here, we focus on the results most relevant to our work and defer a more comprehensive survey to Appendix A. A recent study by (Han et al., 2024) takes initial steps in this direction by analyzing the optimization dynamics of a two-layer CNN trained with diffusion loss on binary class data. Their focus is on contrasting the learning behavior under denoising versus classification objectives, without examining how representation quality evolves across timesteps. In contrast, our work characterizes and compares representations learned at different timesteps, provides a deeper understanding of diffusion-based representation learning and also extends to multi-class settings. A recent study by (Yue et al., 2024) also investigates the influence of timesteps in diffusion-based representation learning, focusing on attribute classification and counterfactual generation. In contrast, our work provides a theoretical explanation and practical metric for the emergence of unimodal representation dynamics and shows its relationship with data and model complexity and training iterations.
@@ 2 Problem Setup
The reverse process transforms noise back into clean data by leveraging the score function $\nabla\log p_{t}(\bm{x}_{t})$ and is governed by the reverse SDE (Anderson, 1982):
@@
Note that the score function $\nabla\log p_{t}(\bm{x}_{t})$ depends on the unknown data distribution $p_{\rm data}$. According to Tweedie’s formula Efron (2011), i.e.,
@@
we can alternatively estimate $\nabla\log p_{t}(\bm{x}_{t})$ by training a network $\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)$, which is referred to as denoising autoencoder (DAE), to approximate the posterior mean $\operatorname{\mathbb{E}}[\bm{x}_{0}|\bm{x}_{t}]$ (Chen et al., 2024b; Xiang et al., 2023; Kadkhodaie et al., 2024). To learn the network parameter $\bm{\theta}$, we minimize the following empirical loss:
@@ 3 Good Distribution Learning Implies Unimodal Representation Dynamics
Building on the setup introduced in Section 2, this section provides a mathematical framework for analyzing the representation quality of diffusion models, and theoretically we characterize the unimodal representation dynamics exhibited by diffusion models across noise levels. We validate our theoretical findings on both synthetic and real-world datasets.
[…]
@@ 3.1 Assumptions of data distribution
In this work, we assume that the underlying data distribution is low-dimensional, which follows a noisy version of the mixture of low-rank Gaussians (MoLRG) (Wang et al., 2024; Elhamifar and Vidal, 2013; Wang et al., 2022), defined as follows.
@@
As shown in Figure 2, data from MoLRG resides on a union of low-dimensional subspaces, each following a Gaussian distribution with a low-rank covariance matrix representing its basis perturbed by some noise. For simplicity, we assume equal subspace dimensions ($d_{1}=\dots=d_{K}=d$), orthogonal bases (i.e., $\bm{U}_{k}^{\star\top}\bm{U}_{l}^{\star}=\bm{0}$ for $k\neq l$), and uniform mixing weights ($\pi_{1}=\dots=\pi_{K}=1/K$). Additionally, the study of noisy MoLRG  distributions is further motivated by the following facts:
  • MoLRG captures the intrinsic low-dimensionality of image data. Although real-world image datasets are high-dimensional in terms of pixel count and data volume, extensive empirical studies (Gong et al., 2019; Pope et al., 2021; Stanczuk et al., 2022) demonstrated that their intrinsic dimensionality is considerably lower. Additionally, recent work (Huang et al., 2024; Liang et al., 2025) has leveraged the intrinsic low-dimensional structure of real-world data to analyze the convergence guarantees of diffusion model sampling. The MoLRG distribution, which models data in a low-dimensional space with rank $d_{k}\ll n$, effectively captures this property.
  • The latent space of latent diffusion models is approximately Gaussian. State-of-the-art large-scale diffusion models (Peebles and Xie, 2023; Podell et al., 2024) typically employ autoencoders (Kingma, 2013) to project images into a low-dimensional latent space, where a KL penalty encourages the learned latent distribution to approximate standard Gaussians (Rombach et al., 2022). Furthermore, recent studies (Jing et al., 2022; Chen et al., 2024b) show that diffusion models can be trained to leverage the intrinsic subspace structure of real-world data.
@@ 3.2 Network parametrization
To analyze how feature representations evolve across noise levels in diffusion models, we study the following nonlinear network parameterization. Specifically, we parameterize the DAE $\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)$ and its corresponding latent feature representation $\bm{h}_{\bm{\theta}}(\bm{x}_{t},t)$ as follows:
@@
This network architecture can be interpreted as a shallow U-Net (Ronneberger et al., 2015) composed with a blockwise mixture-of-experts (Shazeer et al., 2017) mechanism, or equivalently, a restricted form of self-attention, with the following components:
@@
  • Expert weighting. Each projected latent group is then reweighted by a coefficient $\beta_{l}^{t}$, which depends on the input $\bm{x}_{t}$ and timestep $t$ via a softmax function $w_{l}(\bm{x}_{t},t)$. These coefficients form the block-diagonal matrix $\bm{D}(\bm{x}_{t},t)$, which scales each group independently. This structure can be interpreted as a mixture-of-experts model (Shazeer et al., 2017), where the contribution of each expert is modulated by the input. It may also be viewed as a restricted self-attention mechanism (Vaswani et al., 2017), where attention is applied at the group level rather than individually, yielding the feature representations of the input as in (4).
@@ 3.3 A metric for measuring representation quality
Suppose the data $\bm{x}_{0}$ follows the noisy MoLRG introduced in 1. Without loss of generality, let $k\in[K]$ denote the true class of $\bm{x}_{0}$. For DAE $\hat{\bm{x}}_{\bm{\theta}}$ parameterized as in (4), we define:
@@ 3.4 Main theoretical results
Suppose the data $\bm{x}_{0}$ is drawn from a noisy MoLRG data distribution with $K$-class and noise level $\delta$ introduced in 1. Then the optimal $\{\bm{U}\}$ minimizing the loss (2) is the ground truth basis defined in (3), and the optimal DAE $\hat{\bm{x}}_{\bm{\theta}}^{\star}(\bm{x}_{t},t)$ admits the analytical form:
@@
Since the $\bm{U}^{\star}_{l}$-related component captures low-dimensional class-relevant attributes and the $\widetilde{\bm{U}}^{\star}_{l}$-related component captures small-scale, class-irrelevant attributes, the optimal DAE exhibits a fine-to-coarse transition (Choi et al., 2022; Wang and Vastola, 2023; Kamb and Ganguli, 2024) in its output, where the class-irrelevant attributes are progressively removed as the noise level $\sigma_{t}$ increases. Specifically, $\zeta_{t}=\frac{1}{1+\sigma_{t}^{2}}$ quantifies the reduction rate of ${\bm{U}}^{\star}_{l}$ term while $\xi_{t}=\frac{\delta^{2}}{\delta^{2}+\sigma_{t}^{2}}$ quantifies the reduction rate of $\widetilde{\bm{U}}^{\star}_{l}$ term, as $\sigma_{t}$ grows, $\xi_{t}$ decays much more rapidly than $\zeta_{t}$, indicating that the output retains more class-related coarse information while discarding fine-grained, irrelevant details.
@@
(Informal) Suppose that the data $\bm{x}_{0}$ follows the noisy MoLRG introduced in 1. Then the $\mathrm{SNR}$ of the optimal DAE $\hat{\bm{x}}_{\bm{\theta}}^{\star}$ can be approximated as follows:
@@
We defer the formal statement of Theorem 1 and its proof to Section E.2. In the following, we discuss the implications of our result.
@@
Figure 3: Feature probing accuracy and associated $\mathrm{SNR}$ dynamics in MoLRG data. In panel(a) we plot the probing accuracy and $\mathrm{SNR}$ with the feature obtained from a learned DAE $\bm{\hat{x}_{\theta}}$, both of which exhibit a consistent unimodal pattern. The DAE is trained on a 3-class MoLRG dataset with data dimension $n=50$, subspace dimension $d=5$, and noise scale $\delta=0.2$. Additionally, in panel(b) we include the optimal $\mathrm{SNR}$ calculated from the optimal DAE $\bm{\hat{x}_{\theta}}^{\star}$ and the derived approximation in Theorem 1 as a reference.
Figure 4: Illustration of the interplay between the denoising rate and the class confidence rate. The settings follow Figure 3.
@@
Intuitively, our theorem shows that a unimodal curve is mainly induced by the interplay between the “denoising rate” $\sigma_{t}^{2}/\delta^{2}$ and the positive class confidence rate $h(\hat{w}_{t}^{+},\delta)$ as the noise level $\sigma_{t}$ increases. As observed in Figure 4, the “denoising rate” ($\sigma_{t}^{2}/\delta^{2}$) increases monotonically with $\sigma_{t}$, while the class confidence rate $h(\hat{w}_{t}^{+},\delta)$ monotonically declines. Initially, when $\sigma_{t}$ is small, the class confidence rate remains relatively stable due to its flat slope, and an increasing “denoising rate” improves the $\mathrm{SNR}$. However, as indicated by (7), when $\sigma_{t}$ becomes too large, $h(\hat{w}_{t}^{+},\delta)$ quickly decays and approaches $h(\hat{w}_{t}^{-},\delta)$, leading to a drop in the $\mathrm{SNR}$. It is worth noting that though we are mainly characterizing the unimodal dynamics of an approximation of $\mathrm{SNR}$, in practice it closely mimic the trend of the actual $\mathrm{SNR}$ as in Figure 3(b).
@@
As shown in Figures 3 and 5, our theory derived from the noisy MoLRG  distribution effectively captures real-world phenomena. Specifically, we conduct experiments on both synthetic (i.e., noisy MoLRG) and real-world datasets (i.e., CIFAR and ImageNet) to measure $\mathrm{SNR}(\hat{\bm{x}}_{\bm{\theta}},t)$ as well as the feature probing accuracy. For feature probing, we use features extracted at different timesteps as inputs for linear probing. The results consistently show that $\mathrm{SNR}(\hat{\bm{x}}_{\bm{\theta}},t)$ follows a unimodal pattern across all cases, mirroring the trend observed in feature probing accuracy as the noise scale increases. This alignment provides a formal justification for previous empirical findings (Xiang et al., 2023; Baranchuk et al., 2022; Tang et al., 2023), which have reported a unimodal trajectory in the representation dynamics of diffusion models with increasing noise levels. Detailed experimental setups are provided in Appendix D.
@@
Figure 5: Dynamics of feature accuracy and associated $\mathrm{SNR}$ on CIFAR10 and TinyImageNet. Feature accuracy is plotted alongside $\mathrm{SNR}(\hat{\bm{x}}_{\bm{\theta}},t)$. Feature accuracy is evaluated on the test set, while the empirical $\mathrm{SNR}$ is computed from the training set. Both exhibit an aligning unimodal pattern. We use released EDM models (Karras et al., 2022) trained on the CIFAR10 (Krizhevsky et al., 2009) and ImageNet (Deng et al., 2009) datasets, evaluating them on CIFAR10 and TinyImageNet (Li Fei-Fei, 2015), respectively. To compute $\mathrm{SNR}$ , we apply PCA on the CIFAR10/TinyImageNet features to extract the basis $\bm{U}_{l}$s. Further details can be found in Appendix D.

### v3

@@ Abstract
Diffusion models, though originally designed for generative tasks, have demonstrated impressive self-supervised representation learning capabilities. A particularly intriguing phenomenon in these models is the emergence of unimodal representation dynamics, where the quality of learned features peaks at an intermediate noise level. In this work, we conduct a comprehensive theoretical and empirical investigation of this phenomenon. Leveraging the inherent low-dimensionality structure of image data, we theoretically demonstrate that the unimodal dynamic emerges when the diffusion model successfully captures the underlying data distribution. The unimodality arises from an interplay between denoising strength and class confidence across noise scales. Empirically, we further show that, in classification tasks, the presence of unimodal dynamics reliably reflects the generalization of the diffusion model: it emerges when the model generates novel images and gradually transitions to a monotonically decreasing curve as the model begins to memorize the training data.
@@ 1 Introduction
Diffusion models, a class of score-based generative models, have achieved great empirical success in various tasks such as image and video generation, speech and audio synthesis, and solving inverse problems (alkhouri2024diffusion; ho2020denoising; rombach2022high; zhang2024improving; bar2024lumiere; ho2022imagen; kong2020hifi; kongdiffwave2021; roich2022pivotal; ruiz2023dreambooth; chen2024exploring; chung2022improving; song2024solving; li2024decoupled; yang2023lossy). In general, these models, consisting of forward and backward processes, learn data distributions by simulating the non-equilibrium thermodynamic diffusion process (sohl2015deep; ho2020denoising; song2020score). Specifically, the forward process progressively adds Gaussian noise to training samples until they are fully corrupted, while the backward process involves learning a score-based model to generate samples from the noisy inputs (hyvarinen2005estimation; song2020score).
Beyond their strong generative capabilities, recent studies have revealed that diffusion models also possess remarkable representation learning capabilities (baranchuk2021label; xiang2023denoising; mukhopadhyay2023diffusion; chen2024deconstructing; tang2023emergent; han2024feature). In particular, the internal feature extractors of trained diffusion models have been shown to serve as powerful self-supervised learners, achieving strong performance on downstream tasks such as classification (xiang2023denoising; mukhopadhyay2023diffusion), semantic segmentation (baranchuk2021label), and image alignment (tang2023emergent). In many cases, these representations match or even surpass those from specialized self-supervised methods, implying that diffusion models have the potential to unify generative and discriminative learning in vision.
Alongside the empirical success, a prevalent phenomenon, which we refer to as the unimodal representation dynamics, has been widely observed in diffusion models for representation learning (baranchuk2021label; xiang2023denoising; tang2023emergent): the quality of learned representations, as measured by downstream task performance, follows a unimodal trend across noise levels. Specifically, the most effective features consistently appear at an intermediate noise level, while performance degrades as inputs become either fully noisy or entirely clean (see Figure˜1). Despite being widely observed, the underlying cause of this unimodal representation dynamic remains poorly understood.
@@
Figure 1: Unimodal representation dynamics in diffusion-based representation learning tasks. This unimodal representation pattern has been previously observed in diffusion-based representation learning tasks; see (baranchuk2021label; xiang2023denoising; tang2023emergent). To verify this, we train diffusion models on various datasets and evaluate downstream performance using noisy images $\bm{x}_{t}$ at different timesteps $t$. In both classification and segmentation tasks, the performance consistently follows a unimodal trend, peaking at intermediate noise levels. In (b), "mIoU" denotes mean Intersection over Union, a standard metric used in segmentation tasks.
@@
In this work, we investigate the emergence of unimodal representation dynamics in diffusion models and its cause. Our analysis characterizes how representation quality varies across noise levels and offers new perspectives on how diffusion models, despite their generative design, can excel at representation learning. Specifically, we conduct a comprehensive theoretical and empirical study of unimodal representation dynamics in diffusion models. Motivated by the well-established observation that natural image data typically reside on a low-dimensional manifold (gong2019intrinsic; pope2021intrinsic; stanczuk2022your), we examine the representation dynamics through how diffusion models learn from noisy mixtures of low-rank Gaussian (MoLRG) distributions. We show that the emergence of unimodal representations is intrinsically linked to the model’s ability to capture the underlying low-dimensional structure of the data. Our key contributions are as follows:
@@
Recent empirical advances in leveraging diffusion models for downstream representation learning have gained significant attention. However, a theoretical understanding of how diffusion models learn representations across different noise levels remains largely unexplored. Here, we focus on the results most relevant to our work and defer a more comprehensive survey to Appendix˜A. A recent study by (han2024feature) takes initial steps in this direction by analyzing the optimization dynamics of a two-layer CNN trained with diffusion loss on binary class data. Their focus is on contrasting the learning behavior under denoising versus classification objectives, without examining how representation quality evolves across timesteps. In contrast, our work characterizes and compares representations learned at different timesteps, provides a deeper understanding of diffusion-based representation learning and also extends to multi-class settings. A recent study by (yue2024exploring) also investigates the influence of timesteps in diffusion-based representation learning, focusing on attribute classification and counterfactual generation. In contrast, our work provides a theoretical explanation and practical metric for the emergence of unimodal representation dynamics and shows its relationship with data and model complexity and training iterations.
@@ 2 Problem Setup
The reverse process transforms noise back into clean data by leveraging the score function $\nabla\log p_{t}(\bm{x}_{t})$ and is governed by the reverse SDE (anderson1982reverse):
@@
Note that the score function $\nabla\log p_{t}(\bm{x}_{t})$ depends on the unknown data distribution $p_{\rm data}$. According to Tweedie’s formula efron2011tweedie, i.e.,
@@
we can alternatively estimate $\nabla\log p_{t}(\bm{x}_{t})$ by training a network $\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)$, which is referred to as denoising autoencoder (DAE), to approximate the posterior mean $\operatorname{\mathbb{E}}[\bm{x}_{0}|\bm{x}_{t}]$ (chen2024deconstructing; xiang2023denoising; kadkhodaie2023generalization). To learn the network parameter $\bm{\theta}$, we minimize the following empirical loss:
@@ 3 Good Distribution Learning Implies Unimodal Representation Dynamics
Building on the setup introduced in Section˜2, this section provides a mathematical framework for analyzing the representation quality of diffusion models, and theoretically we characterize the unimodal representation dynamics exhibited by diffusion models across noise levels. We validate our theoretical findings on both synthetic and real-world datasets.
Figure 2: An illustration of MoLRG with different noise levels. We visualize samples drawn from noisy MoLRG with noise levels $\delta=0.1,\;0.3$ and $K=3$.
@@ 3.1 Assumptions of data distribution
In this work, we assume that the underlying data distribution is low-dimensional, which follows a noisy version of the mixture of low-rank Gaussians (MoLRG) (wang2024diffusion; elhamifar2013sparse; wang2022convergence), defined as follows.
@@
As shown in Figure˜2, data from MoLRG resides on a union of low-dimensional subspaces, each following a Gaussian distribution with a low-rank covariance matrix representing its basis perturbed by some noise. For simplicity, we assume equal subspace dimensions ($d_{1}=\dots=d_{K}=d$), orthogonal bases (i.e., $\bm{U}_{k}^{\star\top}\bm{U}_{l}^{\star}=\bm{0}$ for $k\neq l$), and uniform mixing weights ($\pi_{1}=\dots=\pi_{K}=1/K$). [^1] The assumptions of subspace orthogonality, equal subspace rank and uniform mixing weights are adopted for analytical simplicity. Empirical results in Appendix C demonstrate that the unimodal dynamics persist when these assumptions are relaxed. Additionally, the study of noisy MoLRG  distributions is further motivated by the following facts:
  • MoLRG captures the intrinsic low-dimensionality of image data. Although real-world image datasets are high-dimensional in terms of pixel count and data volume, extensive empirical studies (gong2019intrinsic; pope2021intrinsic; stanczuk2022your) demonstrated that their intrinsic dimensionality is considerably lower. Additionally, recent work (huang2024denoising; liang2024low) has leveraged the intrinsic low-dimensional structure of real-world data to analyze the convergence guarantees of diffusion model sampling. The MoLRG distribution, which models data in a low-dimensional space with rank $d_{k}\ll n$, effectively captures this property.
  • Latent diffusion models encourage the latent space toward a Gaussian distribution. State-of-the-art large-scale diffusion models (peebles2023scalable; podell2023sdxl) typically employ autoencoders (kingma2013auto) to project images into a latent space, where a KL penalty encourages the learned latent distribution to approximate standard Gaussians (rombach2022high). Furthermore, recent studies (jing2022subspace; chen2024deconstructing) show that diffusion models can be trained to leverage the intrinsic subspace structure of real-world data.
@@
In other words, our data model assumes that the overall dataset lies in a union of low-dimensional subspaces, where each class is associated with a dominant class-specific subspace, and there exists a shared subspace common across all classes that captures class-independent fine-grained details. This shared–specific decomposition has been empirically supported in the context of subspace clustering and representation learning (bousmalis2016domain; zhou2019dual).
@@ 3.2 Network parametrization
To analyze how feature representations evolve across noise levels in diffusion models, we study the following nonlinear network parameterization inspired from the MoLRG data assumption. Specifically, we parameterize the DAE $\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)$ and its corresponding latent feature representation $\bm{h}_{\bm{\theta}}(\bm{x}_{t},t)$ as follows:
@@
This network architecture can be interpreted as a shallow U-Net (ronneberger2015u) composed with a blockwise mixture-of-experts (shazeer2017outrageously) mechanism, or equivalently, a restricted form of self-attention, with the following components:
@@
  • Expert weighting. Each projected latent group is then reweighted by a coefficient $\beta_{l}^{t}$, which depends on the input $\bm{x}_{t}$ and timestep $t$ via a softmax function $w_{l}(\bm{x}_{t},t)$. These coefficients form the block-diagonal matrix $\bm{D}(\bm{x}_{t},t)$, which scales each group independently. This structure can be interpreted as a mixture-of-experts model (shazeer2017outrageously), where the contribution of each expert is modulated by the input. It may also be viewed as a restricted self-attention mechanism (vaswani2017attention), where attention is applied at the group level rather than individually, yielding the feature representations of the input as in (4).
@@ 3.3 A metric for measuring representation quality
Suppose the data $\bm{x}_{0}$ follows the noisy MoLRG introduced in ˜1. Without loss of generality, let $k\in[K]$ denote the true class of $\bm{x}_{0}$. For any trained DAE $\hat{\bm{x}}_{\bm{\theta}}$ parameterized as in (4), we define:
@@ 3.4 Main theoretical results
Suppose the data $\bm{x}_{0}$ is drawn from a noisy MoLRG data distribution with $K$-class and noise level $\delta$ introduced in ˜1. Then the optimal $\{\bm{U}\}$ minimizing the loss (2) is the ground truth basis defined in (3), and the optimal DAE $\hat{\bm{x}}_{\bm{\theta}}^{\star}(\bm{x}_{t},t)$ admits the analytical form:
@@
Since the $\bm{U}^{\star}_{l}$-related component captures low-dimensional class-relevant attributes and the $\widetilde{\bm{U}}^{\star}_{l}$-related component captures small-scale, class-irrelevant attributes, the optimal DAE exhibits a fine-to-coarse transition (choi2022perception; wang2023diffusion; kamb2024analytic) in its output, where the class-irrelevant attributes are progressively removed as the noise level $\sigma_{t}$ increases. Specifically, $\zeta_{t}=\frac{1}{1+\sigma_{t}^{2}}$ quantifies the reduction rate of ${\bm{U}}^{\star}_{l}$ term while $\xi_{t}=\frac{\delta^{2}}{\delta^{2}+\sigma_{t}^{2}}$ quantifies the reduction rate of $\widetilde{\bm{U}}^{\star}_{l}$ term, as $\sigma_{t}$ grows, $\xi_{t}$ decays much more rapidly than $\zeta_{t}$, indicating that the output retains more class-related coarse information while discarding fine-grained, irrelevant details.
@@
(Informal) Suppose that the data $\bm{x}_{0}$ follows the noisy MoLRG introduced in ˜1. Then the $\mathrm{SNR}$ of the optimal DAE $\hat{\bm{x}}_{\bm{\theta}}^{\star}$ can be approximated as follows:
@@
We defer the formal statement of Theorem˜1 and its proof to Section˜E.2. In the following, we discuss the implications of our result.
@@
Figure 3: Feature probing accuracy and associated $\mathrm{SNR}$ dynamics in MoLRG data. In panel(a) we plot the probing accuracy and $\mathrm{SNR}$ with the feature obtained from a learned DAE $\bm{\hat{x}_{\theta}}$, both of which exhibit a consistent unimodal pattern. The DAE is trained on a 3-class MoLRG dataset with data dimension $n=50$, subspace dimension $d=5$, and noise scale $\delta=0.2$. Additionally, in panel(b) we include the optimal $\mathrm{SNR}$ calculated from the optimal DAE $\bm{\hat{x}_{\theta}}^{\star}$ and the derived approximation in Theorem˜1 as a reference.
Figure 4: Illustration of the interplay between the denoising rate and the class confidence rate. The settings follow Figure˜3.
@@
Intuitively, our theorem shows that a unimodal curve is mainly induced by the interplay between the “denoising rate” $\sigma_{t}^{2}/\delta^{2}$ and the positive class confidence rate $h(\hat{w}_{t}^{+},\delta)$ as the noise level $\sigma_{t}$ increases. As observed in Figure˜4, the “denoising rate” ($\sigma_{t}^{2}/\delta^{2}$) increases monotonically with $\sigma_{t}$, while the class confidence rate $h(\hat{w}_{t}^{+},\delta)$ monotonically declines. Initially, when $\sigma_{t}$ is small, the class confidence rate remains relatively stable due to its flat slope, and an increasing “denoising rate” improves the $\mathrm{SNR}$. However, as indicated by (7), when $\sigma_{t}$ becomes too large, $h(\hat{w}_{t}^{+},\delta)$ quickly decays and approaches $h(\hat{w}_{t}^{-},\delta)$, leading to a drop in the $\mathrm{SNR}$. It is worth noting that though we are mainly characterizing the unimodal dynamics of an approximation of $\mathrm{SNR}$, in practice it closely mimic the trend of the actual $\mathrm{SNR}$ as in Figure˜3(b).
@@
As shown in Figures˜3 and 5, our theory derived from the noisy MoLRG  distribution effectively captures real-world phenomena. Specifically, we conduct experiments on both synthetic (i.e., noisy MoLRG) and real-world datasets (i.e., CIFAR and ImageNet) to measure $\mathrm{SNR}(\hat{\bm{x}}_{\bm{\theta}},t)$ as well as the feature probing accuracy. For feature probing, we use features extracted at different timesteps as inputs for linear probing. The results consistently show that $\mathrm{SNR}(\hat{\bm{x}}_{\bm{\theta}},t)$ follows a unimodal pattern across all cases, mirroring the trend observed in feature probing accuracy as the noise scale increases. This alignment provides a formal justification for previous empirical findings (xiang2023denoising; baranchuk2021label; tang2023emergent), which have reported a unimodal trajectory in the representation dynamics of diffusion models with increasing noise levels. Detailed experimental setups are provided in Appendix˜D.
@@
Figure 5: Dynamics of feature accuracy and associated $\mathrm{SNR}$ on CIFAR10 and TinyImageNet. Feature accuracy is plotted alongside $\mathrm{SNR}(\hat{\bm{x}}_{\bm{\theta}},t)$. Feature accuracy is evaluated on the test set, while the empirical $\mathrm{SNR}$ is computed from the training set. Both exhibit an aligning unimodal pattern. We use released EDM models (karras2022elucidating) trained on the CIFAR10 (krizhevsky2009learning) and ImageNet (deng2009imagenet) datasets, evaluating them on CIFAR10 and TinyImageNet (cs231n), respectively. To compute $\mathrm{SNR}$ , we apply PCA on the CIFAR10/TinyImageNet features to extract the basis $\bm{U}_{l}$s. Further details can be found in Appendix˜D.